OpenStudy (anonymous):

How do you write the equation of the line passing through the points (3,-4) and (2,6)?

OpenStudy (anonymous):

finding gradient m=-10 y=mx+c taking any points eg (3,-4) -4=-10(3)+c c=28 therefore y=-10x +28

OpenStudy (anonymous):

formula for gradient m=(y2-y1)/(x2-x1) taake the given points and substitude them on the equation

OpenStudy (anonymous):

How did you get 28?

OpenStudy (aravindg):

plz giv medal

OpenStudy (anonymous):

y=-10x+c sub the give points -4=-10(3)+c c=26 sorry for that of 28

OpenStudy (anonymous):

sorry.made a mistake y=m(x1-x2)+y2 6=m(2-3)-4 10=2m-3m 10=-m m=-10 y=mx+c 6=(-10)2+c 6=-20+c c=26 y=mx+c y=-10x+26